# On the multiplier spectrum of polynomials

### v1

@@ 4.2. Description of the non-injective locus
By conjugating by Möbius transformations, we may assume that $\tau_{1},\tau_{0}$, and $\tilde{h}_{1}$ are polynomials. Set $k_{1}=\deg(\tau_{0})\geq 2$.
We claim that, after further conjugation by a Möbius transformation, the polynomial $\tau_{0}$ has the form $\tau_{0}(z)=z^{k_{1}}$. If $R$ is exceptional, then by [FG22, Theorem 3.39], $h_{1}=\phi_{1}(t)^{\circ N}$ and $\phi_{1}(t)$ are also exceptional, contradicting our assumption. Thus $R$ is a non-exceptional polynomial which is a generalized Lattès map. Since $\tau_{0}:\mathcal{O}_{1}^{\tau_{0}}\to\mathcal{O}_{2}^{\tau_{0}}$ is a covering map between orbifolds, the Riemann–Hurwitz formula gives
$\chi(\mathcal{O}_{2}^{\tau_{0}})=\frac{\chi(\mathcal{O}_{1}^{\tau_{0}})}{k_{1}}<2$
because $k_{1}\geq 2$. In particular, the ramification function of $\mathcal{O}_{2}^{\tau_{0}}$ is not identically $1$. By Lemma 2.4, we have $\nu(\mathcal{O}_{2}^{\tau_{0}})=\{n,n\}$ for some integer $n\geq 2$. By definition of $\mathcal{O}_{2}^{\tau_{0}}$ and considering the point $\infty$, we see $k_{1}=\deg(\tau_{0})\in\nu(\mathcal{O}_{2}^{\tau_{0}})$, hence $n=k_{1}$. Since $\nu(\mathcal{O}_{2}^{\tau_{0}})=\{k_{1},k_{1}\}$, the polynomial $\tau_{0}$ has exactly two critical values in $\widehat{\mathbb{C}}$, hence is of the form $\tau_{0}(z)=az^{k_{1}}+c$, where $a\in\mathbb{C}^{\times}$ and $c\in\mathbb{C}$. Replacing $(\tau_{0},R,\pi)$ by $(L\circ\tau_{0},L\circ R\circ L^{-1},L\circ\pi)$, where L(z)=(zc)/aL(z)=(z-c)/a, we may assume $\tau_{0}(z)=z^{k_{1}}$.
By Theorem 2.8 and Remark 2.9, there exists $R_{1}(z)\in\mathbb{C}[z]$ such that
$R(z)=z^{r_{1}}\cdot R_{1}(z)^{k_{1}}\quad\text{and}\quad\tilde{h}_{1}(z)=z^{r_{1}}\cdot R_{1}(z^{k_{1}}),$
for some integer $r_{1}\geq 1$ with $\gcd(r_{1},k_{1})=1$.
Similarly, if $h_{2}\not\sim R$, there exist $R_{2}(z),\tilde{h}_{2}\in\mathbb{C}[z]$ such that
$h_{2}\sim\tilde{h}_{2},\quad R(z)=z^{r_{2}}\cdot R_{2}(z)^{k_{2}},\quad\tilde{h}_{2}(z)=z^{r_{2}}\cdot R_{2}(z^{k_{2}}),$
for some integers $r_{2}\geq 1$ and $k_{2}\geq 2$ with $\gcd(r_{2},k_{2})=1$. If $h_{2}\sim R$, set $\tilde{h}_{2}=R_{2}=R$, $r_{2}=0$, and $k_{2}=1$.
Set $k={\rm lcm}(k_{1},k_{2})\geq k_{1}\geq 2$ and $k_{j}^{\prime}=k/k_{j}$ for $j=1,2$. Then
@@
Write $R_{j}(z)=z^{l_{j}}V_{j}(z)$ with $l_{j}\in\mathbb{Z}_{\geq 0}$ and $V_{j}(z)\in\mathbb{C}[z]$ satisfying $V_{j}(0)\neq 0$, for $j=1,2$. Then (4.5) becomes
$z^{(r_{1}+k_{1}l_{1})k}\cdot V_{1}(z^{k}))^{k_{1}}=z^{(r_{2}+k_{2}l_{2})k}\cdot V_{2}(z^{k})^{k_{2}}.$
Comparing orders at 0, we get $r_{1}+k_{1}l_{1}=r_{2}+k_{2}l_{2}=:l$. Hence $V_{1}(z^{k})^{k_{1}}=V_{2}(z^{k})^{k_{2}}$, so $V_{1}(z)^{k_{1}}=V_{2}(z)^{k_{2}}$. Examining irreducible factors of $V_{1}(z)$, there exists $V(z)\in\mathbb{C}[z]$ such that $V_{1}(z)=V(z)^{k_{1}^{\prime}}$. Then $V_{2}(z)^{k_{2}}=V(z)^{k}$, which implies $V_{2}(z)=\zeta\cdot V(z)^{k_{2}^{\prime}}$ for some $k_{2}$-th root of unity $\zeta\in\mathbb{C}^{\times}$. Therefore
@@
Similarly, $\gcd(d,k_{2})=1$. Hence $\gcd(d,k_{1}k_{2})=\gcd(d,k)=1$. If $V(z)$ is a constant $c\in\mathbb{C}^{\times}$, then $l=d^{N}\geq 2$ and $h_{1}(z)\sim c^{k_{1}^{\prime}}z^{l}\sim z^{l}$, so all multipliers of $h_{1}$ are in $\mathbb{Z}$. By [JXZ25, Theorem 1.5], $h_{1}=\phi_{1}(t)^{\circ N}$ is exceptional and so is $\phi_{1}(t)$, contradicting our assumption. Thus $V(z)$ is non-constant. In particular,
$d^{N}=\deg(h_{1})=l+k\deg(V)\geq k\geq 2.$
After a suitable iterate, we can make further reductions. Set $V_{1}(z)=V(z)$. For $n\geq 1$, define $V_{n+1}(z)=V(z)^{l^{n}}\cdot V_{n}(z^{l}V(z)^{k})$ inductively. Then $V_{n}(z)\in\mathbb{C}[z]\setminus\mathbb{C}$ and $V_{n}(0)\neq 0$ for all $n\geq 1$. By induction, it is clear that for every $n\geq 1$,
(4.6) $\tilde{h}_{1}^{\circ n}(z)=z^{l^{n}}\cdot V_{n}(z^{k_{1}})^{k_{1}^{\prime}}\quad\text{and}\quad\tilde{h}_{2}^{\circ n}(z)=\zeta^{1+\cdots+l^{n-1}}\cdot z^{l^{n}}\cdot V_{n}(z^{k_{2}})^{k_{2}^{\prime}}.$
Since $\gcd(d,k)=1$ and $l=d^{N}-k\deg(V)$, $\gcd(l,k)=1$. By Euler’s theorem, for the integer $n=k_{2}\varphi(k)\geq 1$, we have
$l^{n}\equiv 1\pmod{k}\quad\text{and}\quad 1+\cdots+l^{n-1}\equiv 0\pmod{k_{2}}$.
@@
$l^{n_{0}}\equiv 1\pmod{k}\quad\text{and}\quad 1+\cdots+l^{n_{0}-1}\equiv 0\pmod{k_{2}}$.
Replacing $(h_{i},\tilde{h}_{i},N,l,V,\zeta)$ by $(h_{i}^{\circ n_{0}},\tilde{h}_{i}^{\circ n_{0}},n_{0}N,l^{n_{0}},V_{n_{0}},1)$ ($i=1,2$), we may assume $l\equiv 1\pmod{k}$ and $\zeta=1$. We still assume $h_{1}\not\sim h_{2}$ and hence $\tilde{h}_{1}\not\sim\tilde{h}_{2}$. In particular, $k_{1}\neq k_{2}$. This concludes the proof. ∎
@@
Pakovich [Pak17, Theorem 1.6] (see also [FG22, Theorem 3.44]) showed that for every polynomial $P$ of degree $d\geq 2$, there exist a polynomial $P_{\rm min}$ of degree $d$ and a non-constant polynomial $\pi_{\rm min}$ with $P\geq_{\pi_{\rm min}}P_{\rm min}$ satisfying the following universal property: for any polynomial $Q\leq P$, there exist polynomials $\pi,\omega$ such that $P\geq_{\omega}Q\geq_{\pi}P_{\rm min}$ and $\pi_{\rm min}=\pi\circ\omega$. Furthermore, $\deg(\pi_{\rm min})$ has an upper bound depending only on $d$.
From the proof above, we immediately see that in Theorem 1.9 (2), the polynomials $(\phi_{1}(t)^{\circ N})_{\rm min}$ and $(\phi_{2}(t)^{\circ N})_{\rm min}$ are mutually semi-conjugate, hence
$(\phi_{1}(t)^{\circ N})_{\rm min}\sim(\phi_{2}(t)^{\circ N})_{\rm min}$
@@
By [FG22, Theorem B], the positive integer $N$ in (1) or (2) has an upper bound depending on $\mathbb{Q}(\phi_{1},\phi_{2})$ and $d$, as $n_{0}\leq k_{2}\varphi(k)$.
@@
Considering a non-empty affine subset of $C$, by Theorem 1.9, there exists a finite subset $S\subset C(\mathbb{C})$ such that for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})\setminus S$, $\phi_{1}(t):=f_{t}$ and $\phi_{2}(t):=g_{t}$ satisfy (1) or (2) of Theorem 1.9. By [FG22, 4. and 5. of Theorem 3.39, 3. of Proposition 3.41], $f_{t}$ and $g_{t}$ are intertwined if they satisfy Theorem 1.9 (1). Note that for all $Q\in\mathbb{C}[z]\setminus\{0\}$ and $m,n\geq 0$, the polynomials $z^{m}Q(z^{n})$ and $z^{m}Q(z)^{n}$ are intertwined [FG22, Theorem 3.39]. Hence by [FG22, Theorem 3.39], we conclude that $f_{t}$ and $g_{t}$ are intertwined if they satisfy Theorem 1.9 (2). Consequently, for all but finitely many $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined. Ji and Xie [JX25, Remark 1.6] showed that the intertwined locus is Zariski-closed in $\mathcal{M}_{d}\times\mathcal{M}_{d}$, hence also Zariski-closed in ${\rm MPoly}^{d}\times{\rm MPoly}^{d}$. Therefore, for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined. ∎
4.3. Ritt move
Case 2 of Theorem 1.9 involves polynomials arising from the “Ritt move” in the Ritt theory of polynomial decompositions [Rit22]; see [ZM08] and [FG22, §3.5.1]. Up to composing with linear polynomials, the only solutions of the equation $P\circ Q=\hat{P}\circ\hat{Q}$ in indecomposable $P,Q,\hat{P},\hat{Q}\in\mathbb{C}[z]$ (of degree $\geq 2$) are the trivial one $P\circ Q=P\circ Q$ and the nontrivial ones has the forms:
@@ 4.3. Ritt move
With additional hypotheses, we can still deduce that the multiplier spectra coincide on an (infinite) arithmetic progression in case (2) of Theorem 1.9, as follows:
@@
According to the proof of Theorem 4.6, we can take $c_{1}$ to be the minimal positive integer such that $\gcd(r(r^{c_{1}}-1),k)=1$ and $d$ to be the minimal positive integer with $r^{d}\equiv 1\pmod{k}$.
@@
If $k=1$, then $P=Q$ and the conclusion is trivial. Assume $k\geq 2$. Write $R(z)=z^{l}\cdot R_{0}(z)$, where $l\in\mathbb{Z}_{\geq 0}$ and $R_{0}(z)\in\mathbb{C}[z]$ with $R_{0}(0)\neq 0$. For $r_{0}=r+lk$ and $M\in\mathbb{Z}_{>0}$ such that $\gcd(r(r^{M}-1),k)=1$, we also have $\gcd(k,r_{0})=\gcd(r_{0}^{M}-1,k)=1$. Replacing $r$ and $R$ by $r_{0}$ and $R_{0}$, respectively, we may assume $R(0)\neq 0$. If $\deg(R)=0$, then both $P$ and $Q$ are conjugate to $z^{r}$ (recall $r\geq 2$), and the conclusion is immediate. Henceforth assume $\deg(R)=s\geq 1$.
First, for every polynomial $f(z)\in\mathbb{C}[z]$ of degree $m\geq 2$ and every $n\in\mathbb{Z}_{>0}$, the point $\infty$ has multiplicity $1$ in ${\rm Fix}(f^{\circ n})$ with $\rho_{f^{\circ n}}(\infty)=0$. Thus we only need to consider periodic points in $\mathbb{C}$. Let ${\rm Fix}(\cdot)$ denote ${\rm Fix}(\cdot,\mathbb{C})$ from now on.
@@
it is easy to see that the map (between multi-sets)
$h:{\rm Fix}(P^{\circ n})\to{\rm Fix}(Q^{\circ n}),\ z_{0}\mapsto z_{0}^{k}$
is well-defined.
Set $F_{1}={\rm Fix}(P^{\circ n})$ and $F_{2}={\rm Fix}(Q^{\circ n})$. For a multi-set $X$ and an element $x\in X$, let ${\rm multi}_{X}(x)$ denote the multiplicity of $x$ in $X$. We claim that for every $z_{0}\in F_{1}$, we have
(4.10) ${\rm multi}_{F_{1}}(z_{0})={\rm multi}_{F_{2}}(z_{0}^{k})\quad\text{and}\quad\rho_{P^{\circ n}}(z_{0})=\rho_{Q^{\circ n}}(z_{0}^{k}).$
Recall that for every rational map $f\in\mathbb{C}(z)$ of degree at least $2$ and a fixed point $z_{0}\in{\rm Fix}(f)$, we have $\rho_{f}(z_{0})\neq 1$ if and only if ${\rm multi}_{{\rm Fix}(f)}(z_{0})=1$. Differentiating (4.9) and evaluating at $z_{0}\in{\rm Fix}(P^{\circ n})$, we obtain
@@
Let $z_{0}\in{\rm Fix}(P^{\circ n})\setminus\{0\}$. By (4.11),

### v2

@@ 4.2. Description of the non-injective locus
After conjugacy by Möbius transformations if necessary, we may assume that $\tau_{1}$, $\tau_{0}$, and $\tilde{h}_{1}$ are polynomials.
We show that, up to conjugacy by a degree one polynomial, we may assume $\tau_{0}(z)=z^{k_{1}}$. If $R$ is exceptional, then by [FG22, Theorem 3.39], $h_{1}=\phi_{1}(t)^{\circ N}$ and $\phi_{1}(t)$ are also exceptional, contradicting our assumption. Thus $R$ is a non-exceptional polynomial that is also a generalized Lattès map. Since $\tau_{0}:\mathcal{O}_{1}^{\tau_{0}}\to\mathcal{O}_{2}^{\tau_{0}}$ is a covering map between orbifolds, the Riemann–Hurwitz formula gives
$\chi(\mathcal{O}_{2}^{\tau_{0}})=\frac{\chi(\mathcal{O}_{1}^{\tau_{0}})}{k_{1}}\leq\frac{2}{k_{1}}\leq 1<2.$
In particular, the ramification function of $\mathcal{O}_{2}^{\tau_{0}}$ is not identically $1$. By Lemma 2.4, we have $\nu(\mathcal{O}_{2}^{\tau_{0}})=\{n,n\}$ for some integer $n\geq 2$. By the definition of $\mathcal{O}_{2}^{\tau_{0}}$, the point $\infty$ shows $k_{1}=\deg(\tau_{0})\in\nu(\mathcal{O}_{2}^{\tau_{0}})$, hence $n=k_{1}$. Since $\nu(\mathcal{O}_{2}^{\tau_{0}})=\{k_{1},k_{1}\}$, the polynomial $\tau_{0}$ has exactly one critical value in $\mathbb{C}$, hence is of the form $\tau_{0}(z)=az^{k_{1}}+c$, where $a\in\mathbb{C}^{\times}$ and $c\in\mathbb{C}$. Replacing the triple $(\tau_{0},R,\tau)$ with $(L\circ\tau_{0},L\circ R\circ L^{-1},L\circ\tau)$, where L(z)=(zc)/aL(z)=(z-c)/a, we may assume $\tau_{0}(z)=z^{k_{1}}$.
By Theorem 2.8 and Remark 2.9, there exists $R_{1}(z)\in\mathbb{C}[z]\setminus\{0\}$ such that
$R(z)=z^{r_{1}}R_{1}(z)^{k_{1}}\quad\text{and}\quad\tilde{h}_{1}(z)=z^{r_{1}}R_{1}(z^{k_{1}}),$
for some integer $r_{1}\geq 1$ with $\gcd(r_{1},k_{1})=1$. Since $R$ is non-exceptional, we have $\deg(R_{1})\geq 1$.
Similarly, if $h_{2}\not\sim R$, then there exist $R_{2}(z),\tilde{h}_{2}\in\mathbb{C}[z]\setminus\mathbb{C}$ such that
$h_{2}\sim\tilde{h}_{2},\quad R(z)=z^{r_{2}}R_{2}(z)^{k_{2}},\quad\tilde{h}_{2}(z)=z^{r_{2}}R_{2}(z^{k_{2}}),$
for some integers $r_{2}\geq 1$ and $k_{2}\geq 2$ with $\gcd(r_{2},k_{2})=1$. If $h_{2}\sim R$, then we set $\tilde{h}_{2}=R_{2}=R$, $r_{2}=0$, and $k_{2}=1$.
Set $k=\mathrm{lcm}(k_{1},k_{2})\geq k_{1}\geq 2$ and $k_{j}^{\prime}=k/k_{j}$ for $1\leq j\leq 2$. Then
@@
Write $R_{j}(z)=z^{l_{j}}V_{j}(z)$ with $l_{j}\in\mathbb{Z}_{\geq 0}$ and $V_{j}(z)\in\mathbb{C}[z]$ satisfying $V_{j}(0)\neq 0$, for $1\leq j\leq 2$. Then (4.8) becomes
$z^{(r_{1}+k_{1}l_{1})k}\cdot V_{1}(z^{k})^{k_{1}}=z^{(r_{2}+k_{2}l_{2})k}\cdot V_{2}(z^{k})^{k_{2}}.$
Comparing the order at 0, we get $r_{1}+k_{1}l_{1}=r_{2}+k_{2}l_{2}=:l\in\mathbb{Z}_{>0}$. We have $V_{1}(z^{k})^{k_{1}}=V_{2}(z^{k})^{k_{2}}$, so $V_{1}(z)^{k_{1}}=V_{2}(z)^{k_{2}}$. Examining irreducible factors of $V_{1}(z)$, there exists $V(z)\in\mathbb{C}[z]$ such that $V_{1}(z)=V(z)^{k_{1}^{\prime}}$. Then $V_{2}(z)^{k_{2}}=V(z)^{k}=V(z)^{k_{2}^{\prime}k_{2}}$, which implies that $V_{2}(z)=\zeta\cdot V(z)^{k_{2}^{\prime}}$ for some $k_{2}$-th root of unity $\zeta\in\mathbb{C}^{\times}$. Therefore
@@
Similarly, $\gcd(d,k_{2})=1$. Then $\gcd(d,k)=1$. Since $R(z)=z^{l}V(z)^{k}$ is non-exceptional, we have $\deg(V)\geq 1$. In particular,
$d^{N}=\deg(R)=l+k\deg(V)\geq r_{1}+k\cdot 1\geq 3.$
The proof of the Claim is completed.
After a further iterate, we can make further reductions. Set $W_{1}(z)=V(z)$. For $n\geq 1$, define $W_{n+1}(z)=W_{n}(z)^{l}\cdot V(z^{l^{n}}W_{n}(z)^{k})$ inductively. Then $W_{n}(z)\in\mathbb{C}[z]\setminus\mathbb{C}$ and $W_{n}(0)\neq 0$ for all $n\geq 1$. By induction, it is clear that for every $n\geq 1$,
$\tilde{h}_{1}^{\circ n}(z)=z^{l^{n}}\cdot W_{n}(z^{k_{1}})^{k_{1}^{\prime}}\quad\text{and}\quad\tilde{h}_{2}^{\circ n}(z)=\zeta^{1+\cdots+l^{n-1}}\cdot z^{l^{n}}\cdot W_{n}(z^{k_{2}})^{k_{2}^{\prime}}$
Since $\gcd(d,k)=1$ and $l=d^{N}-k\deg(V)$, we have $\gcd(l,k)=1$. By Euler’s theorem, for the integer $k_{2}\varphi(k)\in\mathbb{Z}_{>0}$, we have
$l^{k_{2}\varphi(k)}\equiv 1\ (\mathrm{mod}\ k)\quad\text{and}\quad 1+\cdots+l^{k_{2}\varphi(k)-1}\equiv 0\ (\mathrm{mod}\ k_{2}).$
@@
$l^{n_{0}}\equiv 1\ (\mathrm{mod}\ k)\quad\text{and}\quad 1+\cdots+l^{n_{0}-1}\equiv 0\ (\mathrm{mod}\ k_{2}).$
Replacing $(h_{i},\tilde{h}_{i},N,l,V,\zeta)$ by $(h_{i}^{\circ n_{0}},\tilde{h}_{i}^{\circ n_{0}},n_{0}N,l^{n_{0}},W_{n_{0}},1)$ ($i=1,2$), we may assume $l\equiv 1\ (\mathrm{mod}\ k)$ and $\zeta=1$. We still assume $h_{1}\not\sim h_{2}$ and hence $\tilde{h}_{1}\not\sim\tilde{h}_{2}$. In particular, $\tilde{h}_{1}\neq\tilde{h}_{2}$, so $k_{1}\neq k_{2}$. This concludes the proof. ∎
@@
Pakovich [Pak17, Theorem 1.6] (see also [FG22, Theorem 3.44]) showed that for every polynomial $P\in\mathbb{C}[z]$ of degree $d\geq 2$, there exist $\pi_{\mathrm{min}}\in\mathbb{C}[z]\setminus\mathbb{C}$ and a polynomial $P_{\mathrm{min}}\in\mathbb{C}[z]$ of degree $d$ with $P\geq_{\pi_{\mathrm{min}}}P_{\mathrm{min}}$ satisfying the following universal property: for any polynomial $Q\in\mathbb{C}[z]$ with $P\geq Q$, there exist $\pi,\omega\in\mathbb{C}[z]\setminus\mathbb{C}$ such that $P\geq_{\omega}Q\geq_{\pi}P_{\mathrm{min}}$ and $\pi_{\mathrm{min}}=\pi\circ\omega$. Furthermore, $\deg(\pi_{\mathrm{min}})$ is bounded by a constant depending only on $d$.
From the proof above, it follows immediately that in Theorem 1.10 (2), the polynomials $(\phi_{1}(t)^{\circ N})_{\mathrm{min}}$ and $(\phi_{2}(t)^{\circ N})_{\mathrm{min}}$ are mutually semi-conjugate, hence
$(\phi_{1}(t)^{\circ N})_{\mathrm{min}}\sim(\phi_{2}(t)^{\circ N})_{\mathrm{min}}$
@@
Let $K$ be a number field such that $\phi_{1}$, $\phi_{2}$, and $C$ are defined over $K$. By [FG22, Remark 5.21], the positive integer $N$ in (1) or (2) of Theorem 1.10 is bounded by a constant depending only on $K:\mathbb{Q}]$ and $d$, as $n_{0}\leq k_{2}\varphi(k)$.
@@
By restricting to a non-empty affine subset $C_{0}\subseteq C$, we can define a section
$\psi:C_{0}\to\mathrm{Poly}^{d}\times\mathrm{Poly}^{d}$
of
$\Psi\times\Psi:\mathrm{Poly}^{d}\times\mathrm{Poly}^{d}\to\mathrm{MPoly}^{d}\times\mathrm{MPoly}^{d}$
over $C_{0}$. By Theorem 1.10, there exists a finite subset $S\subset C_{0}(\mathbb{C})$ such that for every $t\in C_{0}(\mathbb{C})\setminus S$, $\phi_{1}(t):=\pi_{1}(\psi(t))$ and $\phi_{2}(t):=\pi_{2}(\psi(t))$ satisfy (1) or (2) of Theorem 1.10, where $\pi_{j}:\mathrm{Poly}^{d}\times\mathrm{Poly}^{d}\to\mathrm{Poly}^{d}$ is the $j$-th projection for $1\leq j\leq 2$.
By [FG22, Theorem 3.39(4)(5) and Proposition 3.41(3)], $f_{t}$ and $g_{t}$ are intertwined if they satisfy Theorem 1.10 (1). Note that for all $Q\in\mathbb{C}[z]\setminus\{0\}$ and integers $m,n\geq 0$, the polynomials $z^{m}Q(z^{n})$ and $z^{m}Q(z)^{n}$ are intertwined [FG22, Theorem 3.39]. Again by [FG22, Theorem 3.39], we conclude that $f_{t}$ and $g_{t}$ are intertwined if they satisfy Theorem 1.10 (2). As $S\cup(C\setminus C_{0})$ is finite, for all but finitely many $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined. Ji and Xie [JX25, Remark 1.6] showed that the intertwined locus is Zariski-closed in $\mathcal{M}_{d}\times\mathcal{M}_{d}$, hence also Zariski-closed in $\mathrm{MPoly}^{d}\times\mathrm{MPoly}^{d}$. Therefore, for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined. ∎
4.3. Ritt moves
[…]
@@ 4.3. Ritt moves
With additional hypotheses, we can still deduce that the multiplier spectra coincide over an (infinite) arithmetic progression in case (2) of Theorem 1.10, as follows:
@@
The conclusion trivially holds if $k=1$. When $k\geq 2$, according to the proof of Theorem 4.6, we can take $c_{1}$ to be the minimal positive integer such that $\gcd(r(r^{c_{1}}-1),k)=1$ and $d$ to be the minimal positive integer such that $r^{d}\equiv 1\ (\mathrm{mod}\ k)$.
@@
If $k=1$, then $P=Q$ and the conclusion is trivial. Assume $k\geq 2$. Write $R(z)=z^{l}R_{0}(z)$, where $l\in\mathbb{Z}_{\geq 0}$ and $R_{0}(z)\in\mathbb{C}[z]$ with $R_{0}(0)\neq 0$. Set $r_{0}=r+lk$. For $M\in\mathbb{Z}_{>0}$ such that $\gcd(r(r^{M}-1),k)=1$, we also have $\gcd(r_{0}(r_{0}^{M}-1),k)=1$. Replacing $(r,R)$ by $(r_{0},R_{0})$, we may assume $R(0)\neq 0$. If $\deg(R)=0$, then both $P$ and $Q$ are conjugate to $z^{r}$, and the conclusion is immediate. Henceforth assume $s:=\deg(R)\geq 1$.
Recall that for every rational map $f\in\mathbb{C}(z)$ of degree at least $2$ and a fixed point $z_{0}\in\mathrm{Fix}(f)$, we have $\rho_{f}(z_{0})\neq 1$ if and only if the multiplicity of $z_{0}$ in $\mathrm{Fix}(f)$ is $1$. For every polynomial $f(z)\in\mathbb{C}[z]$ of degree $m\geq 2$ and every $n\in\mathbb{Z}_{>0}$, the point $\infty$ always has multiplier $\rho_{f^{\circ n}}(\infty)=0$, hence $\infty$ has multiplicity $1$ in $\mathrm{Fix}(f^{\circ n})$. Thus for polynomials, we only need to consider periodic points in $\mathbb{C}$. Let $\mathrm{Fix}(\cdot)$ denote $\mathrm{Fix}(\cdot,\mathbb{C})$ from now on.
@@
it is easy to see that the map
$h:\mathrm{Fix}(P^{\circ n})\to\mathrm{Fix}(Q^{\circ n}),\quad z_{0}\mapsto z_{0}^{k}$
is well-defined (as a map between sets).
Set $F_{1}=\mathrm{Fix}(P^{\circ n})$ and $F_{2}=\mathrm{Fix}(Q^{\circ n})$, which are multisets of cardinality $m^{n}$, where $4\leq m:=r+sk=\deg(P)=\deg(Q)$. For a multiset $X$ and an element $x\in X$, let $\mathrm{multi}_{X}(x)$ denote the multiplicity of $x$ in $X$. We claim that for every $z_{0}\in F_{1}$, we have
(4.12) $\mathrm{multi}_{F_{1}}(z_{0})=\mathrm{multi}_{F_{2}}(z_{0}^{k})\quad\text{and}\quad\rho_{P^{\circ n}}(z_{0})=\rho_{Q^{\circ n}}(z_{0}^{k}).$
Differentiating (4.11) and evaluating at $z_{0}\in\mathrm{Fix}(P^{\circ n})$, we obtain
@@
Let $z_{0}\in\mathrm{Fix}(P^{\circ n})$ such that $z_{0}\neq 0$. By (4.13),